OpenStudy (cheddarskittles):

Show work? 3t + 2= 5(t + 6)+2t

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

\[3t+2=5(t+6)+2t\]First distribute the 5.\[3t+2=5t+30+2t\]Then combine like terms.\[3t+2=7t+30\]Combine like terms by adding or subtracting to get the term on the other side of the equal sign.\[3t-7t=30-2\]\[-4t=28\]Finally divide by -4.\[t=-7\]